OpenStudy (anonymous):

2. Show all of your work. Determine whether the pair of lines is parallel, perpendicular, or neither. 2x + 3y = 12 4x+ 6y = –18

OpenStudy (amistre64):

ax +by = c m = -a/b

OpenStudy (amistre64):

-2/3 = -4/6 -2/3 = -2/3 true; parallel

OpenStudy (anonymous):

-2/3=-4/6 -2/3=-2/3: parrallel

OpenStudy (amistre64):

lol ..... its like I have my own parrot :)

OpenStudy (anonymous):

?